OpenStudy (cutiecomittee123):

Which one of these functions does not have a vertical asymptote? Please explain why!!!!!!!! y=5x/x+x^2 y=x/1-x^2 y=5x/1-2x^2 y=5x-1/3+x^2

OpenStudy (bibby):

for vertical asymptotes set the denominator = 0

OpenStudy (cutiecomittee123):

So then it would be x+x^2=0 1-x^2=0 1-2x^2=0 x^2+3=0 Now I am having a brain fart and have no idea how to solve any of these.....

OpenStudy (bibby):

\(x+x^2=x(x+1)=0\\x=0,-1\\...\\1-x^2=0\\1=x^2\\\sqrt1=x\\x=1 \\...\\1-2x^2=0\\1=2x^2\\x^2=\dfrac{1}{2}\)

OpenStudy (bibby):

we're trying to find the denominator with no real roots

OpenStudy (cutiecomittee123):

Oh yeah wow! So wait this is going to sound dumb, but is a fraction a real number?

OpenStudy (bibby):

that's not dumb :) assuming the fraction is rational, yeah http://www.regentsprep.org/regents/math/algebra/AOP1/Lrat.htm

OpenStudy (bibby):

the square root of a negative number isn't real btw

OpenStudy (cutiecomittee123):

yeah thats an imaginary/complex number, i know that

OpenStudy (cutiecomittee123):

So since after multiplying I get x=sqrt-x x=sqrt of 1 x= sqrt of 1/2 and x=sqrt of -3 Would the answer be the last one since it is an irrational number?

OpenStudy (cutiecomittee123):

since the square root of -3 isnt a real number??

OpenStudy (bibby):

yeah :D http://www.wolframalpha.com/input/?i=y%3D%285x-1%29%2F%283%2Bx%5E2%29 this says the domain is all \(\mathbb{R}\) too

OpenStudy (cutiecomittee123):

sweet thanks